Two parallel metal plates carry opposite. Charges
Answer #1

a) E = V/d = 400/(55*10^-3) = 7272.72 V = 7.272 kV


b) Force F = q*E = 2.5*10^-9*7272.72 = 1.818*10^-5 N


C) Work done is W = q*dV = 2.5*10^-9*400 = 1*10^-6 J

d) W = -dU

change in potential energy is dU = -W = - 1*10^-6 J